OpenStudy (anonymous):

Question Navigator 4. (1 pt) Which property justifies this statement? 2 + (8 + 9) = (2 + 8) + 9 A. transitive B. commutative C. distributive D. associative

OpenStudy (plasmataco):

it's not c

OpenStudy (aihberkhan):

Okay this is not really a calculation problem. You simply have to know the meaning of each property. Look at this: Commutative property: When two numbers are added, the sum is the same regardless of the order of the addends. For example 4 + 2 = 2 + 4 Associative Property: When three or more numbers are added, the sum is the same regardless of the grouping of the addends. For example (2 + 3) + 4 = 2 + (3 + 4) Additive Identity Property: The sum of any number and zero is the original number. For example 5 + 0 = 5. Distributive property: The sum of two numbers times a third number is equal to the sum of each addend times the third number. For example 4 * (6 + 3) = 4*6 + 4*3

OpenStudy (aihberkhan):

Now what do you think it is?

OpenStudy (aihberkhan):

Also this may help: http://www.aaamath.com/pro74ax2.htm

OpenStudy (anonymous):

b

OpenStudy (plasmataco):

nice

OpenStudy (aihberkhan):

No... close.. Look at what I just posted.. The definition of each. Try again.

OpenStudy (anonymous):

so its b

OpenStudy (plasmataco):

wait nvm

OpenStudy (aihberkhan):

Nope. Try again.

OpenStudy (plasmataco):

it's not b

OpenStudy (anonymous):

a

OpenStudy (aihberkhan):

Wait.... it is actually associative property! D! Because look at my example: Associative Property: When three or more numbers are added, the sum is the same regardless of the grouping of the addends. For example (2 + 3) + 4 = 2 + (3 + 4)

OpenStudy (plasmataco):

look for the parentheses...

OpenStudy (aihberkhan):

Great Job! :) @owl29

OpenStudy (aihberkhan):

Do you see how I got that? :)

OpenStudy (anonymous):

yes